An amount of P dollars is borrowed for the given length of time at an annual interest rate of r. Find the simple interest thaPlease answer all!!
Answer #1

1)
P = $ 3000, R = 4%, T = 3 months = 3/12 = 0.25 years
SI = PRT/100 = 3000*4*0.25/100 = $ 30

2)
P = $ 5300, R = 2%, T = 7 years
SI = PRT/100 = 5300*2*7/100 = $ 742

3)
F = $ 4004
P = $ 2600
T = 6 years
SI = F - P = $4004 - 2600 = $ 1404
using the formula of SI,
=> SI = PRT/100
=> 1404 = 2600*R*6/100
=> R = 1404*100/ (2600*6)
=> R = 9%

please upvote